A glass maufacturer makes two types of glass . he has 1800 kg silica to make
both type of glass. He uses 216kg more slica to make one type of glass than other type. Find the quantity of slica used in each type.

Answer by Mathtut(3670)

let x and y be the two kinds of silica
:
x+y=1800........eq 1
x=y+216.........eq 2
:
take x's value from eq 2 and plug it into eq 1
:
y+216+y=1800
:
2y=1584
:
y=792
:
so x=792+216=1008
:
x=1008 and y=792